OpenStudy (anonymous):

Find the derivative of f(x) = 2/x at x = -2.

OpenStudy (anonymous):

-1/2

OpenStudy (anonymous):

\[f \prime(x) = \frac{ -2 }{ x^2 }\] then plug in x...

OpenStudy (mathmale):

@Jesi: Please provide guidance to the person asking the question so that that person may find his or her own answers. This is an overriding rule in OpenStudy. Do not just provide answers with no involvement on the part of the person asking the question.

OpenStudy (anonymous):

lol. So when they are set up in fractions you could always find it by adding ^2 to x?

OpenStudy (anonymous):

sorry...

OpenStudy (anonymous):

@yesy7 you can use the idea of negative exponents... :)

OpenStudy (anonymous):

Oh, okay. Thank you!!

OpenStudy (mathmale):

"So when they are set up in fractions you could always find it by adding ^2 to x?" No. You need to know how to apply the Power Rule for Differentiation in this case. You'd find this easier if you'd rewrite the given function as \[f(x)=2x ^{-1}\] and then apply the power rule to that. Can you do that? What is \[\frac{ d }{ dx }x ^{-1}?\]

OpenStudy (mathmale):

Please note: You must know the Power Rule (along with a good number of other rules for differentiation):\[\frac{ d }{ dx }x^n=nx ^{n-1}\]

OpenStudy (mathmale):

In\[\frac{ d }{ dx }x ^{-1},~n=-1\]